# Propaganda and Information Dissemination in the Russo-Ukrainian War : Natural Language Processing of Russian and Western Twitter Narratives

### v1

@@ 2.1.1 Western Tweets
Figure 4 shows the models results and key themes. Our interpretation here reveals a strong focus on the geopolitical and military aspects of the Ukraine conflict. Frequently occurring themes include "Ukrainian Support," "Russian Invasion," "Russiann Disinformation," "European Alliance," and "NATO," which align with theories of geopolitical narratives, emphasising national security and international order [27, 34]. Place-specific keywords such as "Belarus," "Mariupol," and "Kyiv" also highlight the strategic significance of these regions and their role in Ukraine’s resistance and sovereignty [29, 33]. The frequent mention of leaders like "Putin" and "Zelensky" reflects the personalisation of the conflict, a common technique in media representations of war, where individual decisions have significant international ramifications [24].
@@ 2.1.2 Russian Propaganda Tweets
In contrast, Russian propaganda tweets show a markedly different set of themes and narrative strategies. The use of German words like "eingriff," "erweiterung," and "reichstag" indicates efforts to influence German-speaking audiences, reflecting a transnational propaganda strategy. The presence of terms such as "biolabs," "reichstag," and "wokist" also points to the use of conspiracy theories and cultural war rhetoric, aiming to evoke fear and justify military actions [7, 25].
The frequent personalisation of the conflict, through mentions of figures like "Joseph Stalin," "Ali Khamenei," and "Adolf Hitler," is designed to humanise the geopolitical struggle. This approach, known in propaganda studies as the "great man" technique, leverages the symbolic power of leadership figures to influence public opinion [19]. Additionally, the use of derogatory language like "beta," "simp," and "autist" indicates a strategy of trolling and memes in attempt to polarise and degrade people, which is well-documented in studies of online radicalization [10, 11].
@@ 2.2 Sentiment Analysis and Narrative Strategies
Sentiment analysis revealed substantial differences between Russian and Western tweets. Russian propaganda is characterised by a predominantly negative sentiment, using emotionally charged language to provoke fear and justify aggressive actions. Terms like "Nazi," "attack," and "force" frequently appeared, supporting a narrative of existential threat [27, 19]. In contrast, Western tweets exhibited more varied sentiment, often leaning towards neutral or positive tones, especially when discussing humanitarian efforts. This reflects a narrative strategy focused on factual reporting and the ethical obligations to support Ukraine [24].
@@
Our findings show the stark differences in how the conflict is portrayed on Twitter/X, demonstrating the platform’s role as an active battleground for information warfare, where narratives are constructed, contested, and propagated [21, 33].

### v2

@@ 2.1.1 Western Tweets
Figure 4 shows the models results and key themes. Our interpretation here reveals a strong focus on the geopolitical aspects of the Ukraine conflict. Frequently occurring themes include "Ukrainian Support," "Russian Invasion," "Russian Disinformation," "European Alliance," and "NATO," which align with theories of geopolitical narratives, emphasising national security and international order. Keywords such as "Belarus," "Mariupol," and "Kyiv" also highlight the strategic significance of these regions and their role in Ukraine’s resistance and sovereignty. The frequent mention of leaders like "Putin" and "Zelensky" reflects the personalisation of the conflict.
@@ 2.1.2 Russian Propaganda Tweets
In contrast, Russian propaganda tweets show a markedly different set of themes and narrative strategies. The use of German words like "eingriff," "erweiterung," and "reichstag" indicates efforts to influence German-speaking audiences, reflecting a transnational propaganda strategy. The presence of terms such as "biolabs," "reichstag," and "wokist" also points to the use of conspiracy theories and cultural war rhetoric, aiming to evoke fear and justify military actions.
The frequent personalisation of the conflict, through mentions of figures like "Joseph Stalin," "Ali Khamenei," and "Adolf Hitler," is designed to humanise the geopolitical struggle. This approach, known in propaganda studies as the "great man" technique, leverages the symbolic power of leadership figures to influence public opinion [16]. Additionally, the use of derogatory language like "beta," "simp," and "autist" indicates a strategy of trolling and memes in attempt to polarise and degrade people, which is well-documented in studies of online radicalisation.
@@ 2.2 Sentiment Analysis and Narrative Strategies
Sentiment analysis revealed substantial differences between Russian and Western tweets. Russian propaganda is characterised by a predominantly negative sentiment, using emotionally charged language to provoke fear and justify aggressive actions. Terms like "Nazi," "attack," and "force" frequently appeared, supporting a narrative of existential threat [21, 16]. In contrast, Western tweets exhibited more varied sentiment, often leaning towards neutral or positive tones, especially when discussing humanitarian efforts. This reflects a narrative strategy focused on factual reporting and the ethical obligations to support Ukraine.
@@
Our findings show the stark differences in how the conflict is portrayed on Twitter/X, demonstrating the platform’s role as an active battleground for information warfare, where narratives are constructed, contested, and propagated.